OpenStudy (anonymous):

Find the product. x^4 divided by x^4+5x^3 times (x+5)

OpenStudy (anonymous):

\[ \frac {x^4}{ x^4 + 5 x^3}( x+5)=\frac {x^4}{ x^3(x + 5)}( x+5)=x \]

OpenStudy (anonymous):

Of course if \[ x \ne -5 \]

OpenStudy (anonymous):

and \[ x\ne 0\]